DonaldTrumpofQC:

The midpoint of line AC is M (-1, 3). One endpoint is A (4, -1). Find the coordinate of the other endpoint C.

curries:

were u ever given a formula to solve this? if so, i can probably help.

DonaldTrumpofQC:

@curries wrote:
were u ever given a formula to solve this? if so, i can probably help.
There are some we've learned. idk If it's for this. But theirs E M E, then another one

curries:

are u able to send them?

DonaldTrumpofQC:

@curries wrote:
are u able to send them?
idk

curries:

@donaldtrumpofqc wrote:
@curries wrote:
are u able to send them?
idk
try to type the equations out with the equation setting thing

curries:

i didn't see it..

DonaldTrumpofQC:

\[\left( \frac{x_1+x_2}{2}, \frac{y_1+y_2}{2} \right)\]

curries:

oh, thanks

DonaldTrumpofQC:

I'm confused on what the question is asking

curries:

wait let me redo this i slipped up a little..

curries:

bc we're finding out what the coordinates for c is, i suggest replacing the numbers with their respective variables, like x\[x _{a}\] for example

curries:

i'd also add the midpoint equation into the problem in front of it, like M(xₘ, yₘ) = \[\left( \frac{x_1+x_2}{2}, \frac{y_1+y_2}{2} \right)\]

curries:

once u get that, plug in your numbers that u DO have into the equation

curries:

@curries wrote:
i'd also add the midpoint equation into the problem in front of it, like M(xₘ, yₘ) = \[\left( \frac{x_1+x_2}{2}, \frac{y_1+y_2}{2} \right)\]
fixing this, its \[\left( \frac{x_a+x_c}{2}, \frac{y_a+y_c}{2} \right)\]

curries:

after u plug in the m's and a's, show me what u have

DonaldTrumpofQC:

idk what to even do...

DonaldTrumpofQC:

WAIT

curries:

what dont u understand? i can fix some instructions if u need to

DonaldTrumpofQC:

hang on. I'm getting it...

DonaldTrumpofQC:

I'm at 2*3=-1+y 6=-1+y

DonaldTrumpofQC:

the coordinates are (-6,7)

gelphielvr:

@donaldtrumpofqc wrote:
the coordinates are (-6,7)
si- have you figured out how to solve it yet or did you solve it differently than meant to? If you haven't figured it out yet I might be able to help you for future problems like that

DonaldTrumpofQC:

@gelphielvr wrote:
@donaldtrumpofqc wrote:
the coordinates are (-6,7)
si- have you figured out how to solve it yet or did you solve it differently than meant to? If you haven't figured it out yet I might be able to help you for future problems like that
I just said the answer

gelphielvr:

@donaldtrumpofqc wrote:
@gelphielvr wrote:
@donaldtrumpofqc wrote:
the coordinates are (-6,7)
si- have you figured out how to solve it yet or did you solve it differently than meant to? If you haven't figured it out yet I might be able to help you for future problems like that
I just said the answer
i know I was just asking if you know the correct way to solve it now since you didn't before .

DonaldTrumpofQC:

@gelphielvr wrote:
@donaldtrumpofqc wrote:
@gelphielvr wrote:
@donaldtrumpofqc wrote:
the coordinates are (-6,7)
si- have you figured out how to solve it yet or did you solve it differently than meant to? If you haven't figured it out yet I might be able to help you for future problems like that
I just said the answer
i know I was just asking if you know the correct way to solve it now since you didn't before .
I think so